show me the formula that ur are using

Find (a) the present worth P for the cash flows 105 Year 5100 5100 5100 5100 $100 $160 $160 $160
Answer #1

Present value=(CF1)/(1+r)^1+(CF2)/(1+r)^2+(CF3)/(1+r)^3+(CF4)/(1+r)^4+(CF5)/(1+r)^5+(CF6)/(1+r)^6+(CF7)/(1+r)^7+(CF8)/(1+r)^8
Here, CF1 to CF8 are the cash flows from year 1 to 8 and r refers to the discount rate
Substituting the values, we get;

=100/(1+10%)^1+100/(1+10%)^2+100/(1+10%)^3+100/(1+10%)^4+100/(1+10%)^5+160/(1+14%)^6+160/(1+14%)^7+160/(1+14%)^8

=90.90909091+82.6446281+75.13148009+68.30134554+62.09213231+72.89384763+63.9419716+56.08944878
=572.003945 or 572 (Rounded to the nearest whole number)

Present worth for the cash flows =$572